OpenStudy (anonymous):

A candy company places an average of 20 truffles in a one-pound box. The number of truffles per box never varies from the average by more than 2. a. Write an absolute value inequality to represent this situation, using x to represent the actual number of truffles per box. b. Solve the inequality. Show all work leading to your answer

OpenStudy (anonymous):

same as previous question .can u try?

OpenStudy (anonymous):

I suck at multi-step equation. I just suck at math in general

OpenStudy (anonymous):

For the past several months, Shawn’s time to run a mile has averaged 8.5 minutes. Her time never varies from the average by more than 1.25 minutes. a. Write an absolute value inequality to represent this situation, using x to represent Shawn’s actual time to run a mile. b. Solve the inequality. Show all work leading to your answer.

OpenStudy (anonymous):

What about this one

OpenStudy (anonymous):

x-8.5>=1.25 i think.wait 2 minutes

OpenStudy (anonymous):

is it correct?

OpenStudy (anonymous):

x-20<=2

OpenStudy (anonymous):

7.25<x<9.75

OpenStudy (yanasidlinskiy):

A. |x - 20| ≤ 2 B. There are two possible cases that we can have. Case 1: +2 x - 20 = +2 Add 20 to both sides x = 22 Case 2: -2 x - 20 = -2 Add 20 to both sides x = 18 Since the inequality is ≤ we know the solution is the interval between 18 and 22 18 ≤ x ≤ 22 This is for the first question you needed.

OpenStudy (anonymous):

thanks:) do you know the 2nd question? @YanaSidlinskiy

OpenStudy (yanasidlinskiy):

7.25 ≤ t ≤ 9.75 I *think* that one was a bit harder.

OpenStudy (anonymous):

oh okay.

OpenStudy (anonymous):

For the past several months, Shawn’s time to run a mile has averaged 8.5 minutes. Her time never varies from the average by more than 1.25 minutes. a. Write an absolute value inequality to represent this situation, using x to represent Shawn’s actual time to run a mile. b. Solve the inequality. Show all work leading to your answer.

OpenStudy (anonymous):

7.25<x<9.75

OpenStudy (anonymous):

x-8.5>=1.25

OpenStudy (anonymous):

-1.25+8.5<x-8.5<8.5+1.25

OpenStudy (anonymous):

i solved above

OpenStudy (anonymous):

oh sorry i forgot to post the other last question

OpenStudy (anonymous):

nevermind. i think i can do this one on my own

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

same as u posted first question on equality

OpenStudy (anonymous):

anyother

OpenStudy (anonymous):

2. A candy company places an average of 20 truffles in a one-pound box. The number of truffles per box never varies from the average by more than 2. a. Write an absolute value inequality to represent this situation, using x to represent the actual number of truffles per box. b. Solve the inequality. Show all work leading to your answer.

OpenStudy (anonymous):

i and @YanaSidlinskiy already posted above.see from top.thats why i said dont post question within question

OpenStudy (anonymous):

oh okay sorry

OpenStudy (anonymous):

x-20<=2

OpenStudy (anonymous):

its okay.no problem.it confused u